Compare the similarities and differences in the properties of the zn cu fe elements in their combination with water, hydrochloric acid, and nitric acid. Is this behavior consistent with their relative position in the periodic table?
1
Expert's answer
2015-11-17T08:54:34-0500
Zn + 2 H2O = Zn(OH)2 + H2
Zn + 2 HCl = H2 + ZnCl2
3Zn + 8HNO3 = 3Zn(NO3)2 + 2NO + 4H2O

2 Cu + 2 H2O = 2 CuOH + H2
Cu + 2 HCl = CuCl2 + H2
3 Cu + 8 HNO3 = 3 Cu(NO3)2 + 2 NO + 4 H2O

3Fe+4H2O=Fe3O4+4H2
2 Fe(s) + 6 HCl(aq) = 2 FeCl3(aq) + 3 H2(g)
Fe+6HNO3=Fe(NO3)3+3NO2+3H2O

All the indicated metals in the reaction with water and hydrochloric acid displace hydrogen.
In the reaction with the nitric acid all the indicated metals produce nitrates and lead to the nitrogen oxide formation.
This behavior corresponds to their relative position in the periodic table.